Previous Question: A satellite is said to be in a “geosynchronous” orbit if it appears always to remain over the exact same spot on the rotating Earth. What is the period of this orbit? [enter your answer in hours.]

(a) At what distance (in km) from the center of the Earth must such a satellite be placed into orbit?

[Enter your answer in regular decimal notation instead of scientific notation, i.e. not 3.97 x 10^5, but 397000 .]

(b) Why the orbit must be in the plane of Earth’s equator? (choose 1)

- The orbit need not remain in the equatorial plane to be geosynchronous.

- Orbits stay in a plane that includes the earth's center, so an inclined orbit (out of the equatorial plane) will not stay above a fixed point of the surface.

- Because most satellites are launched from locations near the equator, most orbits are in the equatorial plane.

(c) What wavelength of blackbody radiation is emitted from a printed piece of paper with greatest intensity? (I.e., assume the paper is sitting in front of you, you are indoors, and the lights are off. ) Express your answer in microns.

(d) To what region of the EM spectrum does this wavelength correspond? (choose 1)

- microwave

- visible

- ultraviolet

- radio

- infrared
Answer #1

Time period of geosynchronous orbit is almost 1436 min

a) 42164 km distance (in km) from the center of the Earth must such a satellite be placed into orbit

b) Orbits stay in a plane that includes the earth's center, so an inclined orbit (out of the equatorial plane) will not stay above a fixed point of the surface.

c) weins displacement law Tλ = ​b

b = 2898×10−6 m-K​ T = 300 k at room temperature

λ = 2898×10−6 m-K​ /300 K =0.965 * 10-5 m = 9.655 microns

d) Visible